OpenStudy (anonymous):

There is a parade in town there are Large floats and small floats. Large floats are 30ft, small floats are 15ft and there has to be a 10ft space between each float. The parade must be at least 150ft long, but less than 200ft long. What combinations of small and large floats are possible? (They're 7 combinations) Please help.):

OpenStudy (anonymous):

i think your problem is 150<=n(30+10)+m(15+10) -10 <=200

OpenStudy (anonymous):

160<=40n+25m<=210

OpenStudy (anonymous):

n=0 then m can be m=7 and m=8

OpenStudy (anonymous):

160-40n<= 25m <= 210-40n if n=0 160<= 25m <= 210 m=7 and m=8 if n=1 120 <=25m <= 170 m= 5 and m=6 and so on

OpenStudy (anonymous):

Yeah, I guess so... lol. Sooo how should I write it down.?

OpenStudy (anonymous):

set n=2,3,.. in 160-40n<= 25m <= 210-40n then you can find m for each of n as i illustrate at above

OpenStudy (anonymous):

Ok , that makes sense. Haha, I really apreciate you helping me.(: Thankss.

OpenStudy (anonymous):

im not sure, you can help from other people

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

n=Number of large floats m=Number of small floats i tack it 1) 0 of large floats and 7 of small floats 2) 0 of large floats and 8 of small floats 3) 1 of large floats and 5 of small floats 4) 1 of large floats and 6 of small floats

OpenStudy (anonymous):